Mary has 14 hats. She has 4 green, 2 blue and 8 red. If she selects two hats what is the probability that both hats are different colors?
Round your answer to two decimal places.

Answer by ewatrrr(10682)

 
Hi, Previously posted
Mary has 30 hats. She has 5 green, 7 blue, 12 red, and 6 brown.
selects two hats at random
P(they are different colors) = 1 - P(both are the same color)
P(both are the same color) = 5/30*4/29 + 7/30*6/29 + 12/30*11/29 + 6/30*5*29
= (20 + 42 + 132 + 30)/30*29 = 224/30*29 = .2575
P(they are different colors) = 1 -.2575 = .7425